OpenStudy (anonymous):

5^(x −x^2) = 1/ (625^x) solve the equation for x, smaller value and larger value.... I really don't get this one :(

OpenStudy (freckles):

I know 625 can be written as 25^2 and 25 can be written as 5^2 \[625=25^2=(5^2)^2=5^4\] And you have in your problem\[\frac{1}{625^x}=625^{-x}=(5^4)^{-x}=5^{-4x}\]

OpenStudy (freckles):

hope that helps

OpenStudy (anonymous):

Thank you so much